OpenStudy (anonymous):

Select the inequality that corresponds to the given graph. a number line graph with a closed circle at 4, to the right and a closed circle at negative 1, shading to the left. 5x + 7 greater than or equal to 2 and 10x less than or greater to 30 6x + 7 greater than or equal to 25 or 2x - 8 < -4 4x + 10 less than or greater to 6 or 5x - 6 greater than or equal to 14 37 less than or greater to 8x - 3 less than or greater to 61

OpenStudy (campbell_st):

the closed circle means to include the number shading to the right means with a closed circle is greater than or equal to shading to the left with a closed circle means less than or equal to you are looking for a pair of inequalities with the solution \[x \le -1 ... or.... x \ge 4\] you need to solve the inequalities, just like you would solve an equation Option 1 \[5x + 7 \ge 2\] subtract 7 from both sides \[5x \ge -5\] divide both sides by 5 \[x \ge -1\] you need less than -1 so choice 1 is eliminated. try the rest